OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n? n ≥ - the quantity 4 times c minus 3 all over p n ≥ - the quantity 4 times c minus 13 all over p n ≤ - the quantity 4 times c minus 3 all over p n ≤ - the quantity 4 times c minus 13 all over p

OpenStudy (anonymous):

recursive function is shown below: f(1) = 5 and f(n) = f(n - 1) - 6; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 5, 1, 7, 13, 19, ... 5, -1, -7, -13, -19, ... 5, 11, 17, 23, 29, ... 5, -11, -17, -23, -29, ...

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

just these three problems

OpenStudy (anonymous):

you want the answers or do u want to know how to solve it?

jigglypuff314 (jigglypuff314):

It would be more preferable to only post one question at a time

OpenStudy (anonymous):

its just three i dint think itd be a problem

jigglypuff314 (jigglypuff314):

the equation of the line that passes through the points (-2, 1) and (1, 10)? first find the slope between the two given points, can you do that?

OpenStudy (anonymous):

hardly

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

2-1, 1-10

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

mmm more like 10 - 1 ------ = ? 1 - (-2)

OpenStudy (anonymous):

9/ 3

OpenStudy (anonymous):

PLEASE HELP WITH OTHER TWO

jigglypuff314 (jigglypuff314):

I wasn't done with the first... so now you have 9/3 = 3 as a slope then you have the format y = mx + b where m = slope and b = y-intercept you can plug in 3 = slope = m so y = 3x + b then we are going to temporarily plug in one of the original points as (x, y) to solve for b do you know how to do that? :)

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

maybe if you explain

jigglypuff314 (jigglypuff314):

we solved for the slope and got 3 because 9/3 = 3 then we have a format called the slope-intercept form which is y = mx + b when m = slope and b = y-intercept so since the slope is 3 then m = 3 so plug that in and get y = 3x + b but what is b? well we would then have to temporarily plug in an original point to solve for b like plug in x = -2 and y = 1 so 1 = 3(-2) + b and solve for b from there